the length of a rectangle is 3 feet less than five times its width, if the perimeter is 90 feet, find the dimensions (lenth and Width). What is the area of the rectangle?
Answer by mananth(16946)

width =w
length = 5w-3
Perimeter = 2(L+W)
2(w+5w-3)=90
2(6w-3)=90
/2
6w-3=45
6w=48
/6
w=8
width = 8 feet
length = 5*8-3=37 feet
Area = L*w
Area = 8*37=296 sq.ft